There are two hinged vertical mirrors with an angle of

140 °. A horizontal light beam falls on one of these

mirrors at a certain angle and is then reflected from

both mirrors. What is the acute angle between the

beam incident on the first mirror and the beam

reflected from the second mirror?


1
Expert's answer
2022-03-24T16:55:55-0400

As indicated in the figure,If i be the angle of incidence, r be the angle of reflection,δ=π−(π−2i−2r)Hence, δ=2i+2rAlso, using the sum of angles of a triangle, i+r−θ=0Thus, δ=2θWhich means, that the deviation is independent of either the angle of incidence and also the angle of relection. It only depends on the angle between the mirrors